John is planning to join a gym. Gym A charges a $100 sign-up fee to join plus $30 per month. Gym B has a smaller sign-up fee of $25 but charges $35 per month.
when will the two gyms charge the same amount?

Answer by ikleyn(52834)

.
John is planning to join a gym. Gym A charges a $100 sign-up fee to join plus $30 per month.
Gym B has a smaller sign-up fee of $25 but charges $35 per month.
when will the two gyms charge the same amount?
~~~~~~~~~~~~~~~~~

    100 + 30m = 25 + 35m,


where m is the number of months.


From this equation

    100 - 25 = 35m - 30m

       75    =    5m

        m    =    75/5 = 15.


ANSWER.   The charges sare the same after 15 months.

Solved.